OpenStudy (anonymous):

how to convert (-1,squareroot 3) to polar where r is negative

OpenStudy (anonymous):

\(r\) is a length, so it is not negative

OpenStudy (anonymous):

\[r=\sqrt{a^2+b^2}=\sqrt{1^2+\sqrt{3}^2}=\sqrt{1+3}=\sqrt{4}=2\] \[\theta =\frac{4\pi}{3}\]

OpenStudy (anonymous):

r has to be negative

OpenStudy (anonymous):

so it can flip when graphing it

OpenStudy (anonymous):

then make is \(-2\) and change the sign inside

OpenStudy (anonymous):

the answer is (-2,5pi/3)

OpenStudy (anonymous):

i just dont know how to find the angle once i find r